### v1

@@ 3.2 Mean charged-particle multiplicity
Figure 3: Mean charged-particle multiplicity within QCD jets at the LHC. Left: Inclusive mean charged-particle multiplicity as a function of jet transverse momentum $p_{T}$, comparing ATLAS data ATLAS:2019rqw , PYTHIA simulations, DLA, and N3LO results. Right: Mean charged-particle multiplicity in quark and gluon jets from PYTHIA, DLA, and N3LO predictions. For the DLA results, $K_{\text{LPHD}}=0.80$ is obtained by fitting the inclusive mean multiplicity data shown in the left panel with $Q_{0}=0.5$ GeV held fixed, while for the N3LO results $K=0.0353$ and $Q_{0}=0.053$ GeV are both determined from fits to the same data.
@@
In this subsection, we briefly summarize the results of mean charged-particle multiplicities used in our studies of inclusive multiplicity distributions, as given in eq. (3.1). Specifically, we employ the DLA results with running coupling, given in eqs. (15) and (19), as well as the next-to-next-to-next-to-leading-order (N3LO) results from refs. Dremin:1999ji ; Capella:1999ms , which are summarized in appendix A (see also ref. Dremin:2000ep for a review). To facilitate a meaningful comparison between parton-level theoretical predictions and experimental charged-particle data in ref. ATLAS:2019rqw , we make use of the Local Parton–Hadron Duality (LPHD) hypothesis Azimov:1984np ; Dokshitzer:1995ev . Within this approach, we introduce a normalization factor $K_{\text{LPHD}}$ for the DLA results, defined as the ratio of the mean charged-particle multiplicity to the mean parton multiplicity. The corresponding normalization parameter in the N3LO results is denoted by $K$.
@@
To confront the theoretical calculations for the inclusive mean charged-particle multiplicity with the measurements in ref. ATLAS:2019rqw , we first calculate the fractions of quark and gluon jets at LO using eqs. (40) and (41). Here, we consider two leading jets with $|\eta|<2.1$ over the $p_{T}$ range 0.10.12.52.5 TeV in $pp$ collisions at $\sqrt{s}=13$ TeV. The PDFs are taken from the CT18NLO via LHAPDF Buckley:2014ana . Using these extracted jet fractions, we then fix the parameters in the theoretical calculations by fitting to the experimental data for the inclusive mean charged-particle multiplicity in jets with radius $R=0.4$.
@@
The left panel of figure 3 compares the theoretical results of the inclusive mean multiplicity with ATLAS data and PYTHIA simulations. The DLA mean multiplicities depend on two parameters, $Q_{0}$ and $K_{\text{LPHD}}$. We fix $Q_{0}=0.5$ GeV, to be consistent with the discussions in sec. 2.2.2 and ref. Duan:2025ngi , and determine $K_{\text{LPHD}}$ through a $\chi^{2}$ fit to the ATLAS data ATLAS:2019rqw , obtaining $K_{\text{LPHD}}=0.80$ with $\chi^{2}/\text{d.o.f.}=23.1/13=1.78$.333 We have also tried a $\chi^{2}$ fit to the experimental data to determine both parameters. However, this yields a value of $Q_{0}$ very close to $\Lambda$, due to the infrared divergence in the limit $Q_{0}\to\Lambda$ (see eq. (19)), without improving the $\chi^{2}$. In such fits, the global minimum of $\chi^{2}$ does exist, but it occurs at a highly perturbative value of $Q_{0}>20$ GeV. For the N3LO results, the two parameters $Q_{0}$ and $K$ are both determined via fits to the ATLAS data, yielding $K=0.0353$ and $Q_{0}=0.053$ GeV with $\chi^{2}/\text{d.o.f.}=0.904/12=0.075$. As shown in the plot, the N3LO predictions provide an impressively accurate description of the data, while the DLA results remain reasonable but tend to underestimate $\bar{n}_{ch}$ at low $p_{T}$ and overestimate it at high $p_{T}$. The PYTHIA simulations, with MPI and hadronization included, also reproduce the measurements well, aside from a mild overshoot at high $p_{T}$.
@@ 3.3 Inclusive multiplicity distributions
where we have used the relation $n=x\bar{n}=x_{q}\bar{n}_{q}=x_{g}\bar{n}_{g}$. Here, the fractions $r_{q}$ and $r_{g}$ are determined from the LO cross section, as in the mean multiplicity calculations. For the KNO scaling functions $\Psi_{a}$, we adopt the MDLA results with $\gamma_{0}=0.43$ and $c_{q}=0.8$, as described in sec. 2.2.2, since the DLA predictions clearly fail to describe the measurements reported in ref. ATLAS:2019rqw . The remaining relevant quantity is the ratio $\bar{n}_{g}/\bar{n}_{q}$ appearing in $\bar{n}/\bar{n}_{a}$.
Figure 5: Investigation of KNO scaling via inclusive multiplicity distributions in pppp collisions. Shown here are our theoretical results using the MDLA KNO scaling functions ($\Psi_{\text{MDLA}}$), the LO cross section ($\sigma_{\text{LO}}$), and the ratios of mean multiplicities between gluon and quark jets in DLA ($\bar{n}_{\text{DLA}}$) and N3LO ($\bar{n}_{\text{N${}^{3}$LO}}$), compared with the ATLAS measurements of $\bar{n}_{ch}P(n_{ch})$ reported in ref. ATLAS:2019rqw across eight $p_{T}$ bins.
@@
In figure 5, our theoretical predictions are also compared with the measurements reported in ref. ATLAS:2019rqw . The results obtained with the N3LO mean multiplicities show very good agreement with the experimental data across all reported $p_{T}$ bins. In contrast, predictions using DLA mean multiplicities provide a good description only in the range $0.1~\text{TeV}<p_{T}<0.2~\text{TeV}$, as shown in the top left panel. For the remaining $p_{T}$ bins, clear discrepancies appear, even when accounting for the relatively large uncertainties and coarse binning ($\Delta n_{ch}=4$).
@@
Figure 6: Inclusive multiplicity distributions in pppp collisions. Our theoretical predictions, shown in comparison to the $P(n_{ch})$ measurements reported in ref. ATLAS:2019rqw , are the same as those in figure 5, divided by the corresponding theoretical results of the inclusive mean multiplicity. The figure also displays the PYTHIA results, with a bin width of $\Delta n_{ch}=4$ in each panel.
@@
Figure 6 presents the same theoretical results as in figure 5, now divided by the corresponding theoretical inclusive mean multiplicities and compared directly with the $P(n_{ch})$ measurements from ref. ATLAS:2019rqw . The predictions with N3LO mean multiplicities continue to agree very well with the experimental data across the entire $p_{T}$ range, as the N3LO inclusive mean multiplicity provides a very good description of the ATLAS measurements, as shown in the left panel of figure 3. In contrast, results obtained with DLA mean multiplicities again show less accurate agreement, though with a noticeable improvement for $2.0~\text{TeV}<p_{T}<2.5~\text{TeV}$ (and a deterioration for $0.1~\text{TeV}<p_{T}<0.2~\text{TeV}$). This behavior reflects the deviations of the DLA inclusive mean multiplicity from the data in the low- and high-$p_{T}$ bins.
@@
The same figure also shows comparisons with PYTHIA simulations. The PYTHIA $P(n_{ch})$ distributions are presented using the same bin width of $\Delta n_{ch}=4$ as in the ATLAS measurements. As shown in this figure, our theoretical predictions are consistent with the PYTHIA results across all $p_{T}$ bins. The agreement with PYTHIA is expected, as the values of $\gamma_{0}$ and $c_{q}$ were determined by fitting to PYTHIA results in ref. Duan:2025ngi , as shown in figure 2. In turn, the consistency between our theoretical predictions and experimental measurements validates our approach of fixing $\gamma_{0}$ and $c_{q}$ based on PYTHIA simulations, although a more sophisticated determination of these parameters could be achieved by fitting directly to experimental data.
@@
Thus, the agreement of our theoretical results with the ATLAS measurements provides further evidence that the measurements of ref. ATLAS:2019rqw can be consistently described in terms of two distinct KNO scaling functions, as presented within MDLA in sec. 2.2.2, in line with our earlier observation using PYTHIA in ref. (Duan:2025ngi, ). Nonetheless, our theoretical calculations can be further refined, as it relies on the LO cross section, neglects KNO scaling violations induced by the running coupling (see ref. Dokshitzer:2025owq for a discussion at large $n_{ch}/\bar{n}_{ch}$) and treats $c_{q}$ as a free fitting parameter. A more systematic analysis, which would require improved theoretical inputs for each quantity involved, as discussed at the beginning of this subsection, is deferred to future work.
@@ 4 Multiplicity distributions in quark and gluon jets versus jet topics
As discussed above, the analysis of inclusive multiplicity distributions requires additional theoretical inputs, such as $r_{a}$ and $\bar{n}_{a}$, besides the KNO scaling functions, which makes establishing KNO scaling less straightforward. A more direct investigation of KNO scaling in QCD jets can be carried out alongside quark–gluon discrimination using jet substructure techniques (see refs.Marzani:2019hun ; Larkoski:2024uoc for recent reviews). The feasibility of this approach has been demonstrated with PYTHIA simulations in ref. Duan:2025ngi , where energy–correlation functions (ECFs) Larkoski:2013eya are employed to probe KNO scaling. However, because ECF-based studies select jets within a constrained phase space, the MDLA results presented in sec. 2 are not directly applicable. In this section, we instead compare the MDLA KNO scaling functions for quark and gluon jets with results obtained via the topic modeling method Metodiev:2018ftz ; Komiske:2018vkc applied to the two jet samples measured by ATLAS in ref. ATLAS:2019rqw .
In ref. ATLAS:2019rqw , the two selected leading jets are separated into two samples based on the pseudorapidities of the two jets: the more forward jet sample and the more central jet sample. The charged-particle multiplicity distributions with a bin width of $\Delta n_{ch}=4$ are measured for both jet samples, denoted as $h_{i}^{f}$ and $h_{i}^{c}$, where $i$ labels the $n_{ch}$ bins. According to topic modeling Metodiev:2018ftz ; Komiske:2018vkc , the multiplicity distributions in quark and gluon jets can be approximated by the multiplicity distributions for topics 1 ($T_{1}$) and 2 ($T_{2}$):
@@
Figure 7: Charge-particle multiplicity distributions in quark (topic 1) and gluon (topic 2) jets. The plots compare the MDLA results with the multiplicity distributions for topics 1 and 2 extracted from ATLAS data ATLAS:2019rqw . In the experimental results, error bars represent statistical uncertainties only, while shaded bands present systematic uncertainties. Also included are PYTHIA results, where quark and gluon jets are identified by the underlying hard processes in PYTHIA Duan:2025ngi (dashed histograms) and via the topic modeling approach Metodiev:2018ftz ; Komiske:2018vkc (solid histograms).
Figure 8: Mean charged-particle multiplicities in quark (topic 1) and gluon (topic 2) jets. The figure shows the average multiplicities for topics 1 and 2 extracted by applying topic modeling to the ATLAS more-forward and more-central jet samples (dots; statistical errors shown as error bars, systematic uncertainties as shaded bands), compared with the ATLAS results using the nominal quark and gluon fractions from the PYTHIA simulation (triangles) in ref. ATLAS:2019rqw , with experimental uncertainties omitted, as well as with PYTHIA results from ref. Duan:2025ngi (dotted) and PYTHIA results obtained via topic modeling (solid).
Figure 7 compares the MDLA results, divided by $\bar{n}_{a}$, with the multiplicity distributions for $T_{1}$ and $T_{2}$ extracted from ATLAS data, as well as the corresponding PYTHIA results. Here, $\bar{n}_{q}$ and $\bar{n}_{g}$ are taken as the mean multiplicities for $T_{1}$ and $T_{2}$, respectively. To validate the topic modeling method, we first compare the charged-particle distributions of quark and gluon jets identified through the hard processes in PYTHIA, as implemented in ref. Duan:2025ngi , with those obtained via topic modeling applied to PYTHIA-generated jet samples. Overall, the two results are consistent across all $p_{T}$ bins, though the degree of agreement varies among different $p_{T}$ ranges. For example, in the range $0.3~\mathrm{TeV}<p_{T}<0.4~\mathrm{TeV}$, as shown in the top left panel, when restricting to $n_{ch}$ bins with probability above 0.01 to eliminate those dominated by statistical uncertainties, the gluon jet results agree at the percent level, and for quark jets only a single $n_{ch}$ bin shows a relative deviation exceeding 10%.
@@
As shown in figure 7, the extracted jet-topic distributions from the ATLAS samples are consistent with the PYTHIA results once uncertainties are taken into account. Notably, for $0.9~\mathrm{TeV}<p_{T}<1.0~\mathrm{TeV}$, the extracted distributions from experimental data and PYTHIA, whether topic modeling is used or not, are consistent, differing from the results exemplified in ref. ATLAS:2019rqw , which we attribute to our exclusion of $n_{ch}$ bins with large uncertainties. At higher $p_{T}$ (see the bottom right panel), the agreement between our experimental extractions and the PYTHIA results deteriorates, as the central values of the two ATLAS jet samples become more alike than those in PYTHIA.
As shown in figure 8, the mean multiplicities extracted from ATLAS jet samples via topic modeling differ from those reported in ref. ATLAS:2019rqw using the same method, due to the exclusion of $n_{ch}$ bins with large uncertainties. In contrast, they agree with the results obtained via the alternative method using the nominal quark and gluon fractions from the PYTHIA simulation provided by ATLAS in ref. ATLAS:2019rqw and are consistent with the PYTHIA predictions within propagated uncertainties across the full $p_{T}$ range, although the central values lie below the PYTHIA results at higher $p_{T}$ for gluon jets. Here, the large uncertainties for topics 1 and 2 at both low and high $p_{T}$ arise from the near-similarity of the two jet samples, as mentioned above.
@@
Figure 9: KNO scaling in jet topics. Left: multiplicity distributions extracted from the two ATLAS jet samples ATLAS:2019rqw using topic modeling, with propagated experimental uncertainties (see figure 7) omitted for clarity. Right: corresponding distributions from PYTHIA jet samples. In both panels, the bin width is $n_{ch}=4$, and the results are compared with the MDLA scaling functions (dashed lines).
As summarized in figure 9, the extracted multiplicity distributions for the two jet topics exhibit reasonably distinct scaling behavior. For the PYTHIA results, they are consistent with the distributions extracted in ref. Duan:2025ngi , and the MDLA predictions provide a good description when the slight spread in PYTHIA results is neglected. For the extracted ATLAS results, the central values also display similar scaling, with quark jets (Topic 1) agreeing better with the MDLA results than gluon jets (Topic 2). While caution is required when drawing final conclusions due to experimental uncertainties and coarse binning as explicitly shown in figures 7 and 8, these observations demonstrate that jet topic modeling offers a promising approach for extracting multiplicity distributions and supporting studies of KNO scaling.

### v2

@@ 3.2 Mean charged-particle multiplicity
Figure 3: Mean charged-particle multiplicity within QCD jets at the LHC. Left: Inclusive mean charged-particle multiplicity as a function of jet transverse momentum $p_{T}$, comparing ATLAS data Aad and others (2019), PYTHIA simulations, DLA, and N3LO results. Right: Mean charged-particle multiplicity in quark and gluon jets from PYTHIA, DLA, and N3LO predictions. For the DLA results, $K_{\text{LPHD}}=0.80$ is obtained by fitting the inclusive mean multiplicity data shown in the left panel with $Q_{0}=0.5$ GeV held fixed, while for the N3LO results $K=0.0353$ and $Q_{0}=0.053$ GeV are both determined from fits to the same data.
@@
In this subsection, we briefly summarize the results of mean charged-particle multiplicities used in our studies of inclusive multiplicity distributions, as given in eq. (3.1). Specifically, we employ the DLA results with running coupling, given in eqs. (15) and (19), as well as the next-to-next-to-next-to-leading-order (N3LO) results from refs. Dremin and Gary (1999); Capella et al. (2000), which are summarized in appendix A (see also ref. Dremin and Gary (2001) for a review). To facilitate a meaningful comparison between parton-level theoretical predictions and experimental charged-particle data in ref. Aad and others (2019), we make use of the Local Parton–Hadron Duality (LPHD) hypothesis Azimov et al. (1985); Dokshitzer et al. (1996). Within this approach, we introduce a normalization factor $K_{\text{LPHD}}$ for the DLA results, defined as the ratio of the mean charged-particle multiplicity to the mean parton multiplicity. The corresponding normalization parameter in the N3LO results is denoted by $K$.
@@
To confront the theoretical calculations for the inclusive mean charged-particle multiplicity with the measurements in ref. Aad and others (2019), we first calculate the fractions of quark and gluon jets at LO using eqs. (40) and (41). Here, we consider two leading jets with $|\eta|<2.1$ over the $p_{T}$ range 0.10.12.52.5 TeV in $pp$ collisions at $\sqrt{s}=13$ TeV. The PDFs are taken from the CT18NLO via LHAPDF Buckley et al. (2015). Using these extracted jet fractions, we then fix the parameters in the theoretical calculations by fitting to the experimental data for the inclusive mean charged-particle multiplicity in jets with radius $R=0.4$.
@@
The left panel of figure 3 compares the theoretical results of the inclusive mean multiplicity with ATLAS data and PYTHIA simulations. The DLA mean multiplicities depend on two parameters, $Q_{0}$ and $K_{\text{LPHD}}$. We fix $Q_{0}=0.5$ GeV, to be consistent with the discussions in sec. 2.2.2 and ref. Duan et al. (2025), and determine $K_{\text{LPHD}}$ through a $\chi^{2}$ fit to the ATLAS data Aad and others (2019), obtaining $K_{\text{LPHD}}=0.80$ with $\chi^{2}/\text{d.o.f.}=23.1/13=1.78$. [^3] We have also tried a $\chi^{2}$ fit to the experimental data to determine both parameters. However, this yields a value of $Q_{0}$ very close to $\Lambda$, due to the infrared divergence in the limit $Q_{0}\to\Lambda$ (see eq. (19)), without improving the $\chi^{2}$. In such fits, the global minimum of $\chi^{2}$ does exist, but it occurs at a highly perturbative value of $Q_{0}>20$ GeV. For the N3LO results, the two parameters $Q_{0}$ and $K$ are both determined via fits to the ATLAS data, yielding $K=0.0353$ and $Q_{0}=0.053$ GeV with $\chi^{2}/\text{d.o.f.}=0.904/12=0.075$. As shown in the plot, the N3LO predictions provide an impressively accurate description of the data, while the DLA results remain reasonable but tend to underestimate $\bar{n}_{ch}$ at low $p_{T}$ and overestimate it at high $p_{T}$. The PYTHIA simulations, with MPI and hadronization included, also reproduce the measurements well, aside from a mild overshoot at high $p_{T}$.
@@ 3.3 Inclusive multiplicity distributions
where we have used the relation $n=x\bar{n}=x_{q}\bar{n}_{q}=x_{g}\bar{n}_{g}$. Here, the fractions $r_{q}$ and $r_{g}$ are determined from the LO cross section, as in the mean multiplicity calculations. For the KNO scaling functions $\Psi_{a}$, we adopt the MDLA results with $\gamma_{0}=0.43$ and $c_{q}=0.8$, as described in sec. 2.2.2, since the DLA predictions clearly fail to describe the measurements reported in ref. Aad and others (2019). The remaining relevant quantity is the ratio $\bar{n}_{g}/\bar{n}_{q}$ appearing in $\bar{n}/\bar{n}_{a}$.
Figure 5: Investigation of KNO scaling via inclusive multiplicity distributions in pppp collisions. Shown here are our theoretical results using the MDLA KNO scaling functions ($\Psi_{\text{MDLA}}$), the LO cross section ($\sigma_{\text{LO}}$), and the ratios of mean multiplicities between gluon and quark jets in DLA ($\bar{n}_{\text{DLA}}$) and N3LO ($\bar{n}_{\text{N${}^{3}$LO}}$), compared with the ATLAS measurements of $\bar{n}_{ch}P(n_{ch})$ reported in ref. Aad and others (2019) across eight $p_{T}$ bins.
@@
In figure 5, our theoretical predictions are also compared with the measurements reported in ref. Aad and others (2019). The results obtained with the N3LO mean multiplicities show very good agreement with the experimental data across all reported $p_{T}$ bins. In contrast, predictions using DLA mean multiplicities provide a good description only in the range $0.1~\text{TeV}<p_{T}<0.2~\text{TeV}$, as shown in the top left panel. For the remaining $p_{T}$ bins, clear discrepancies appear, even when accounting for the relatively large uncertainties and coarse binning ($\Delta n_{ch}=4$).
@@
Figure 6: Inclusive multiplicity distributions in pppp collisions. Our theoretical predictions, shown in comparison to the $P(n_{ch})$ measurements reported in ref. Aad and others (2019), are the same as those in figure 5, divided by the corresponding theoretical results of the inclusive mean multiplicity. The figure also displays the PYTHIA results, with a bin width of $\Delta n_{ch}=4$ in each panel.
@@
Figure 6 presents the same theoretical results as in figure 5, now divided by the corresponding theoretical inclusive mean multiplicities and compared directly with the $P(n_{ch})$ measurements from ref. Aad and others (2019). The predictions with N3LO mean multiplicities continue to agree very well with the experimental data across the entire $p_{T}$ range, as the N3LO inclusive mean multiplicity provides a very good description of the ATLAS measurements, as shown in the left panel of figure 3. In contrast, results obtained with DLA mean multiplicities again show less accurate agreement, though with a noticeable improvement for $2.0~\text{TeV}<p_{T}<2.5~\text{TeV}$ (and a deterioration for $0.1~\text{TeV}<p_{T}<0.2~\text{TeV}$). This behavior reflects the deviations of the DLA inclusive mean multiplicity from the data in the low- and high-$p_{T}$ bins.
@@
The same figure also shows comparisons with PYTHIA simulations. The PYTHIA $P(n_{ch})$ distributions are presented using the same bin width of $\Delta n_{ch}=4$ as in the ATLAS measurements. As shown in this figure, our theoretical predictions are consistent with the PYTHIA results across all $p_{T}$ bins. The agreement with PYTHIA is expected, as the values of $\gamma_{0}$ and $c_{q}$ were determined by fitting to PYTHIA results in ref. Duan et al. (2025), as shown in figure 2. In turn, the consistency between our theoretical predictions and experimental measurements validates our approach of fixing $\gamma_{0}$ and $c_{q}$ based on PYTHIA simulations, although a more sophisticated determination of these parameters could be achieved by fitting directly to experimental data.
@@
Thus, the agreement of our theoretical results with the ATLAS measurements provides further evidence that the measurements of ref. Aad and others (2019) can be consistently described in terms of two distinct KNO scaling functions, as presented within MDLA in sec. 2.2.2, in line with our earlier observation using PYTHIA in ref. (Duan et al., 2025). Nonetheless, our theoretical calculations can be further refined, as it relies on the LO cross section, neglects KNO scaling violations induced by the running coupling (see ref. Dokshitzer and Webber (2025a) for a discussion at large $n_{ch}/\bar{n}_{ch}$) and treats $c_{q}$ as a free fitting parameter. A more systematic analysis, which would require improved theoretical inputs for each quantity involved, as discussed at the beginning of this subsection, is deferred to future work.
@@ 4 Multiplicity distributions in quark and gluon jets versus jet topics
As discussed above, the analysis of inclusive multiplicity distributions requires additional theoretical inputs, such as $r_{a}$ and $\bar{n}_{a}$, besides the KNO scaling functions, which makes establishing KNO scaling less straightforward. A more direct investigation of KNO scaling in QCD jets can be carried out alongside quark–gluon discrimination using jet substructure techniques (see refs.Marzani et al. (2019); Larkoski (2024) for recent reviews). The feasibility of this approach has been demonstrated with PYTHIA simulations in ref. Duan et al. (2025), where energy–correlation functions (ECFs) Larkoski et al. (2013) are employed to probe KNO scaling. However, because ECF-based studies select jets within a constrained phase space, the MDLA results presented in sec. 2 are not directly applicable. In this section, we instead compare the MDLA KNO scaling functions for quark and gluon jets with results obtained via the topic modeling method Metodiev and Thaler (2018); Komiske et al. (2018) applied to the two jet samples measured by ATLAS in ref. Aad and others (2019).
In ref. Aad and others (2019), the two selected leading jets are separated into two samples based on the pseudorapidities of the two jets: the more forward jet sample and the more central jet sample. The charged-particle multiplicity distributions with a bin width of $\Delta n_{ch}=4$ are measured for both jet samples, denoted as $h_{i}^{f}$ and $h_{i}^{c}$, where $i$ labels the $n_{ch}$ bins. According to topic modeling Metodiev and Thaler (2018); Komiske et al. (2018), the multiplicity distributions in quark and gluon jets can be approximated by the multiplicity distributions for topics 1 ($T_{1}$) and 2 ($T_{2}$):
@@
Figure 7: Charge-particle multiplicity distributions in quark (topic 1) and gluon (topic 2) jets. The plots compare the MDLA results (dashed) with the multiplicity distributions for topics 1 and 2 extracted from ATLAS data Aad and others (2019) (dots). In the experimental results, the error bars indicate statistical uncertainties, while the shaded bands correspond to systematic uncertainties. For comparison, predictions from PYTHIA are also shown, where quark and gluon jets are identified according to the underlying hard processes Duan et al. (2025) (dotted) and via the topic modeling approach Metodiev and Thaler (2018); Komiske et al. (2018) (solid).
Figure 8: Mean charged-particle multiplicities in quark (topic 1) and gluon (topic 2) jets. The figure shows the average multiplicities for topics 1 and 2 extracted by applying topic modeling to the ATLAS more-forward and more-central jet samples (dots; statistical errors shown as error bars, systematic uncertainties as shaded bands), compared with the ATLAS results using the nominal quark and gluon fractions from the PYTHIA simulation (triangles) in ref. Aad and others (2019), with experimental uncertainties omitted, as well as with PYTHIA results from ref. Duan et al. (2025) (dotted) and PYTHIA results obtained via topic modeling (solid).
Figure 7 compares the MDLA results, divided by $\bar{n}_{a}$, with the multiplicity distributions for $T_{1}$ and $T_{2}$ extracted from ATLAS data, as well as the corresponding PYTHIA results. Here, $\bar{n}_{q}$ and $\bar{n}_{g}$ are taken as the mean multiplicities for $T_{1}$ and $T_{2}$, respectively. To validate the topic modeling method, we first compare the charged-particle distributions of quark and gluon jets identified through the hard processes in PYTHIA, as implemented in ref. Duan et al. (2025), with those obtained via topic modeling applied to PYTHIA-generated jet samples. Overall, the two results are consistent across all $p_{T}$ bins, though the degree of agreement varies among different $p_{T}$ ranges. For example, in the range $0.3~\mathrm{TeV}<p_{T}<0.4~\mathrm{TeV}$, as shown in the top left panel, when restricting to $n_{ch}$ bins with probability above 0.01 to eliminate those dominated by statistical uncertainties, the gluon jet results agree at the percent level, and for quark jets only a single $n_{ch}$ bin shows a relative deviation exceeding 10%.
@@
As shown in figure 7, the extracted jet-topic distributions from the ATLAS samples are consistent with the PYTHIA results once uncertainties are taken into account. Notably, for $0.9~\mathrm{TeV}<p_{T}<1.0~\mathrm{TeV}$, the extracted distributions from experimental data and PYTHIA, whether topic modeling is used or not, are consistent, differing from the results exemplified in ref. Aad and others (2019), which we attribute to our exclusion of $n_{ch}$ bins with large uncertainties. At higher $p_{T}$ (see the bottom right panel), the agreement between our experimental extractions and the PYTHIA results deteriorates, as the central values of the two ATLAS jet samples become more alike than those in PYTHIA.
As shown in figure 8, the mean multiplicities extracted from ATLAS jet samples via topic modeling differ from those reported in ref. Aad and others (2019) using the same method, due to the exclusion of $n_{ch}$ bins with large uncertainties. In contrast, they agree with the results obtained via the alternative method using the nominal quark and gluon fractions from the PYTHIA simulation provided by ATLAS in ref. Aad and others (2019) and are consistent with the PYTHIA predictions within propagated uncertainties across the full $p_{T}$ range, although the central values lie below the PYTHIA results at higher $p_{T}$ for gluon jets. Here, the large uncertainties for topics 1 and 2 at both low and high $p_{T}$ arise from the near-similarity of the two jet samples, as mentioned above.
@@
Figure 9: KNO scaling in jet topics. Left: Multiplicity distributions extracted from the two ATLAS jet samples Aad and others (2019) using topic modeling, with propagated experimental uncertainties (see figure 7) omitted for clarity. Right: Corresponding distributions from PYTHIA jet samples. In both panels, the bin width is $n_{ch}=4$, and the results are compared with the MDLA scaling functions (dashed).
As summarized in figure 9, the extracted multiplicity distributions for the two jet topics exhibit reasonably distinct scaling behavior. For the PYTHIA results, they are consistent with the distributions extracted in ref. Duan et al. (2025), and the MDLA predictions provide a good description when the slight spread in PYTHIA results is neglected. For the extracted ATLAS results, the central values also display similar scaling, with quark jets (Topic 1) agreeing better with the MDLA results than gluon jets (Topic 2). While caution is required when drawing final conclusions due to experimental uncertainties and coarse binning as explicitly shown in figures 7 and 8, these observations demonstrate that jet topic modeling offers a promising approach for extracting multiplicity distributions and supporting studies of KNO scaling.